The area of the right angled triangle is 600 sg cm. If the base of the triangle exceeds the altitude by 10 cm. find the dimensions of the triangle
 

Grade:

1 Answers

Shanya Singhal
9 years ago
(x)(x+10)  = 600 sq. cm  (area of right angled triangle)
     2
So,
x+ 10x – 1200 = 0
(x - 30)(x + 40) = 0
x = 30 and x = – 40
rejecting x = – 40 as dimentions of a triangle cannot be in negative
Therefore,
ANSWER= Base = x = 30 cm and Altitude = (x + 10) = 40 cm and hypotenuse = 50 cm (By Pythagorus Theorum)